17. Let f: I → R, where I is an open interval containing the point c, and let k ∈ R. Prove the following.

(a) f is differentiable at c with f ′(c) = k iff limh→0 [ f (c + h) – f (c)]/h = k.

*(b) If f is differentiable at c with f ′(c) = k, then limh→ 0 [ f (c + h) – f (c – h)]/2h = k.

(c) If f is differentiable at c with f ′(c) = k, then lim n →∞ n[f (c + 1/n) – f (c)] = k.

(d) Find counterexamples to show that the converses of parts (b) and (c) are not true. 


The book is Steven R. Lay, Analysis with an introduction to proof.


1
Expert's answer
2021-08-02T17:17:19-0400

Answer:-

We have given "f:I \rightarrow R \ \ \ \ ;" I is an open interval is differentiable at "c \in (a,b)"

and "f'(c)=k \ \ \ \ ;k\in IR"

then f(x) can be approximated by linear function of x , "\ \ \ \because" k is a constant

"\therefore" "f(x)=kx +d \ \ \ \ \ \ \ \ d\in IR"

Now "f(c+h)=k(c+h)+d=kc+kh+d"

"f(c)=kc+d"

"\therefore lim_{h\rightarrow 0}\frac{f(c+h)-f(c)}{h}\\"

"= lim_{h\rightarrow 0} \frac{kc+kh+d-kc-d}{h}"

"=lim_{h\rightarrow 0}\frac{kh}{h}\\
=\boxed{k}"

conversely

"lim_{h\rightarrow 0}\frac{f(c+h)-f(c)}{h}=k"

Consider two points on graph of "f(x) , [c+h,f(c+h)] , [c,f(c)]"

now slope of any line passing through the points is "\frac{f(c+h)-f(c)}{(c+h)-(c)}"

As "lim_{h\rightarrow 0 }" "\frac{f(c+h)-f(c)}{(c+h)-(c)}" becomes slope of tangent at point x=c , when is also derivative of function at x=c .

"f'(c)=lim_{\rightarrow 0}\frac{f(c+h)-f(c)}{h}"

"\implies \boxed{f'(c)=k}"


(b)

given , f'(c)=k

"\implies lim_{h\rightarrow 0 }\frac{f(c+h)-f(c)}{h}=k.............(1)\\"

Replace c by c-h

"\implies lim_{h\rightarrow 0 }\frac{f(c-h+h)-f(c-h)}{h}=k.............(1)\\
\implies lim_{h\rightarrow 0 }\frac{f(c)-f(c-h)}{h}=k.............(2)\\"

add (1) and (2)

"\implies\boxed{ lim_{h\rightarrow 0 }\frac{f(c+h)-f(c-h)}{2h}=k}"


(c)

We have given ,

"f'(c)=k"

"\implies lim_{h\rightarrow 0 }\frac{f(c+h)-f(c)}{h}=k"

Replace h by "1\over n"

"\implies As \ h\rightarrow 0 \ \ , n\rightarrow \infin"

"\implies lim_{n\rightarrow 0 }[f(c+\frac{1}{n}-f(c)]=k"

Counter example for converse of (b)


(d)

"f(x) = \begin{cases}
 2x+3 &\text{; }x\ge 2 \\
 x+5 &\text{l; } x 
<2
\end{cases}"


"lim _{h\rightarrow 0} \frac{(2+h)-f(2-h)}{2h}"

"=lim _{h\rightarrow 0} \frac{(2(2+h)+3)-f(2-h)}{2h}"

"=lim _{h\rightarrow 0} \frac{4+2h+3-2+h-5}{2h}"

"= \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{3}{2}=\boxed{k}"

now ,

"lim _{h\rightarrow 0} \frac{(c+h)-f(c)}{h}"

"=lim _{h\rightarrow 0} \frac{(2+h)-f(2)}{h}"

"=lim _{h\rightarrow 0} \frac{2(2+h)+3-f(2(2)+3)}{h}"

"=lim _{h\rightarrow 0}\frac{2h}{h}"

"\boxed{2}\ne k \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ [k=\frac{3}{2}]"